integrate x^2+x+5/(x^2+4)(x+1)dx
1
Expert's answer
2020-06-30T17:28:46-0400

Let 5/((x2+4)(x+1))=(Ax+B)/(x2+4)+C/(x+1).

Then 5=(Ax+B)(x+1)+C(x2+4)

5=(A+C)x2+(B+A)x+(B+4C)

We obtain system

A+C=0,

B+A=0,

B+4C=5.

Then B=-A, C=-A=>-A-4A=5=>-5A=-5=>A=-1,B=1,C=1.

And 5/((x2+4)(x+1))=(-x+1)/(x2+4)+1/(x+1).

So ∫ (x2+x+5/((x2+4)(x+1)))dx="\int" (x2+x+(-x+1)/(x2+4)+1/(x+1))dx=

"\int"(x2+x-x/(x2+4)+1/(x2+4)+1/(x+1))dx=x3/3+x2/2-(ln(x2+4))/2+(arctan(x/2))/2+ln(x+1)+C